Payments of $670 are being made at the end of each month for 5 years at an interest rate of 8% compounded monthly. Calculate the Present value?
1
Expert's answer
2019-11-26T14:18:27-0500
"n=12 \times 5=60"

"r= \frac {0.08}{12}= 0.00667"


"670 \times \frac {1-(1+ 0.00667)^{-60}}{0.00667}=33043.35"

Answer: 33043.35.